# A Notion of Complexity for Theory of Mind via Discrete World Models

### v2

@@ Abstract
Theory of Mind (ToM) can be used to assess the capabilities of Large Language Models (LLMs) in complex scenarios where social reasoning is required. While the research community has proposed many ToM benchmarks, their hardness varies greatly, and their complexity is not well defined. This work proposes a framework to measure the complexity of ToM tasks. We quantify a problem’s complexity as the number of states necessary to solve it correctly. Our complexity measure also accounts for spurious states of a ToM problem designed to make it apparently harder. We use our method to assess the complexity of five widely adopted ToM benchmarks. On top of this framework, we design a prompting technique that augments the information available to a model with a description of how the environment changes with the agents’ interactions. We name this technique Discrete World Models (DWM) and show how it elicits superior performance on ToM tasks.
@@ 1 Introduction
Theory of Mind (ToM) studies how agents form and use beliefs to reason in dynamic environments Premack and Woodruff (1978). Originally developed to describe human interactions Preston and De Waal (2002); Tomasello (2009) as well as toddlers’ psychological development Wimmer and Perner (1983); Baron-Cohen et al. (1985), ToM has been quickly adopted by other fields, including artificial intelligence McCarthy (1979); Scassellati (2002) and machine learning Rabinowitz et al. (2018). In machine learning, ToM has both descriptive and prescriptive usage: on the one hand, ToM benchmarks assess the capabilities of a model in complex environments; on the other, ToM’s frameworks such as theory-theory Gopnik and Wellman (1994) and simulation theory Churchland (2013) have been widely adopted to test the proficiency of Large Language Models (LLMs) in social tasks where humans excel Strachan et al. (2024).
@@
In this work, we propose a framework to characterise a ToM benchmark’s difficulty, i.e., its complexity, as the number of state events that are sufficient to track the state of an object, including $k^{\text{th}}$-order beliefs. We characterise the complexity of five standard ToM benchmarks, from false belief to commonsense and social reasoning, and compute their complexity as a proxy of their inherent difficulty. Inspired by prompting techniques that split a task into elementary sub-problems that are solved sequentially, like Tree of Thoughts Yao et al. (2023) and least-to-most prompting Zhou et al. (2023a), we introduce a technique that stimulates a model’s reasoning capabilities via Discrete World Models (DWM). DWM leverages the notion of statefulness via a succinct and coherent representation of each state events, as illustrated in Figure 1. We test DWM on ToMi Le et al. (2019), MindGames Sileo and Lernould (2023), Adv-CSFB  Shapira et al. (2023), SocialIQA Sap et al. (2019), and FANToM Kim et al. (2023), eliciting superior performance than Chain of Thoughts (CoT) Wei et al. (2022) and Tree of Thoughts (ToT) Yao et al. (2023) on those problems whose state spaces are informative. We further assess whether memorisation affects a model’s performance, and we discover that while this phenomenon happens for standard benchmarks such as ToMi Le et al. (2019), with input-output pairs that can be retrieved word for word via prompting, it does not correlate with a drop of performance. We conduct our experiments on a variety of open- and closed-source LLMs, including GPT-3.5-Turbo, GPT-4 OpenAI (2023), LLaMA3-70B AI@Meta (2024) and Mixtral 8x7B Jiang et al. (2024). In summary, in this paper:
  • We introduce the concept of complexity of a ToM task to quantify the hardness of keeping track of the elements (e.g., agents’ beliefs or objects’ states) that are sufficient to produce the correct answer to different problems.
@@
We consider our work a step towards a framework that formalizes the hardness of a ToM problem univocally inspired by the theory of World Models Wong et al. (2023).
@@ 2 Related Work
Over 40 years of research on ToM in psychology Premack and Woodruff (1978); Baron-Cohen et al. (1985); Dennett (1988); Wellman (2017) on human development has created a fertile ground for the development of these ideas in adjacent fields. In the last decade, many works studied ToM in artificial intelligence and machine learning Baker et al. (2011); Rabinowitz et al. (2018), with applications to multi-agent systems and reinforcement learning Gronauer and Diepold (2022). More recently, the rise in popularity of LLMs shifted the interest towards understanding and benchmarking large models’ capacity to solve increasingly complex ToM tasks Aru et al. (2023); Zhou et al. (2023b); Mahowald et al. (2024). While some researchers believe LLMs have already become proficient in solving ToM tasks Bubeck et al. (2023); Kosinski (2023); Strachan et al. (2024), others show scepticism and illustrate cases where they fail on trivial variations of well-known problems Ullman (2023); Shapira et al. (2023); Sap et al. (2023). In a joint effort between computer scientists and psychologists, many ToM benchmarks have been developed and used to test neural-network models, including LLMs Gandhi et al. (2022); Chen et al. (2024); Strachan et al. (2024). Recently, concepts such as World Models Ha and Schmidhuber (2018) have found applicability and mostly as discrete prompting techniques in conjunction with optimisation procedures Hao et al. (2023); Moghaddam and Honey (2023). Researchers have found evidence of an emergent internal representation (e.g., World Model’s surrogates) of the state games Li et al. (2022); Toshniwal et al. (2021) and state-tracking abilities Li et al. (2021); Kim and Schuster (2023); Kim et al. (2024), necessary for correct belief tracking in ToM problems. Inspired by the work in Zhou et al. (2023a) and the results in Zhou et al. (2023b). Our prompting technique is inspired by Park et al. (2023) and Nye et al. (2021): the former develops an architecture to record the agent’s experiences. The latter proposes a prompting technique that forces a model to express the intermediate computational steps to solve a problem.
Figure 2: How statefulness and statelessness (Def. 3.2) are computed for the motivating example in Fig. 1. For ${obj_{1}}$, an optimal split to track the apple merges the first two states and chunks of the input prompt. For ${obj_{2}}$, which involves the $1^{\text{st}}$-order belief of Bob, the statefulness is higher, with $e_{2}$ that cannot be merged with $e_{3}$ as it introduces partial observability. The complexity of the task (bottom) is computed as per Eq. 2, with the complexity of stateless objects that is discounted as not directly relevant to the question/answer.
@@ 3 Methodology
In this section, we introduce a notion of complexity for ToM problems: such notion quantifies the hardness of a problem as the number of computational steps humans take to solve them. We then present the DWM prompting technique within the complexity framework and show how it differs from standard methods like CoT and ToT. We further characterise its efficiency with the number of input/output tokens and queries to a model as the control variables.
@@ 3.1 On the Complexity of ToM
Providing a consistent representation of the environment, including each agent’s beliefs, inspired us to characterise the complexity of a ToM problem in terms of sufficient elements to track to output the correct result. Consider a problem prompt $p$, expressed in natural language, that describes how multiple agents interact with an environment object $\mathbf{obj}$, as illustrated in Figure 2 (top). In our framework, an object can be the state of the apple as well as the $k^{\text{th}}$-order belief of an agent about the apple position. Our framework naturally extends to multiple objects by considering their union.
Suppose that in $p$, the state of $\mathbf{obj}$ is modified $T>0$ times, thus identifying $T$ unique configurations, namely $E_{\mathbf{obj}}=\{e_{1},..,e_{T}\}$. To correctly solve a ToM task where $p$ is complemented by a query about $\mathbf{obj}$, a model should distinguish between the interactions that modify the configuration of $\mathbf{obj}$, i.e., the stateful states, from those that modify any other stateless object $Obj\setminus\mathbf{obj}$, i.e., those one does not need to track. We first show how to define the cost of tracking a task’s stateful states, which we complement with that of the stateless. Both definitions concur in defining the complexity of a ToM task.
@@ 3.1.1 Stateful and Stateless Complexity
A state event for an object $\mathbf{obj}$ is an event that links adjacent state descriptions that involve, for both the environment state $e_{t}$ and the sub-prompt $p_{\leq t}$, a state change of $\mathbf{obj}$. Formally, we define a relation, $F_{\mathbf{obj}}$, to specify which pairs of state descriptions form a state event: $F_{\mathbf{obj}}((e_{t},p_{\leq t}),(e_{t+1},p_{\leq t+1}))\equiv e_{t}\neq e_{t+1}\ \wedge\ p_{\leq t+1}=p_{\leq t}\oplus p_{t+1}$ where $1\leq t\leq\lvert p\rvert.$ ($\lvert p\rvert$ denotes the number of atomic prompts.)
@@ 3.1.2 The Complexity of a ToM Task
We can now define the notion of statefulness of a ToM task specified as a prompt $p$ as the size of Eq. 3.2, namely $T_{\mathbf{obj}}=\lvert E_{\mathbf{obj}}\rvert$. The process of computing the statefulness of an object or its belief is illustrated in Fig. 2.
For a ToM task where the question to solve relates to an object $\mathbf{obj}$, one must ensure that changes to any other object, namely $Obj\setminus\mathbf{obj}$, do not affect $\mathbf{obj}$. While tracking the evolution of what is irrelevant to answer the question is unnecessary, a computation model must assess whether a particular environmental change affected $\mathbf{obj}$. We thus introduce the notion of statelessness, i.e., the cost of discerning whether a change in the environment affects $\mathbf{obj}$. The computation is similar to that of Def. 3.2, with $\mathbf{obj}$ that is replaced by any object in $Obj\setminus\mathbf{obj}$; yet, for stateless objects, we introduce a discount factor $\tau$ to penalise the complexity of state events that do not affect $\mathbf{obj}$. Mathematically, we formalise the statelessness of a ToM task involving an object $\mathbf{obj}$ as $\tau\sum_{obj\in Obj\setminus\mathbf{obj}}T_{obj}$.
@@ 3.2 Discrete World Models
A (Large) Language Model is a function that predicts the next token (out of a finite vocabulary) conditioned on the sequence of previously fed/generated tokens, namely $\psi:\mathbf{v}\in V^{*}\xrightarrow{}v\in V$. Such a mechanism can be used to sample multiple token outputs until an ‘end-of-text’ token is predicted by invoking $\psi$ in an auto-regressive fashion, i.e., $\psi(v|\mathbf{v})$. In our setting, a problem is specified as a tuple $(p,Q)$, where $p$ is a ToM problem and $Q$ is a query function that modifies $p$ according to a prompting technique, namely $Q:p\xrightarrow{}p^{\prime}$. The LLMs output $y$ for an input $Q(p)$ is then compared for correctness against an oracle $\Omega$, i.e., $\Omega:\psi(Q(p))\xrightarrow{}\{0,1\}$, where $1$ means correct classification (00, otherwise). On a sample of $N>0$ ToM problems, the accuracy of a model $\psi$ is then measured as $\frac{1}{N}\sum_{i=1}^{N}\Omega(\psi(Q(p_{i}))$, i.e., the average number of times a model is correct in its prediction.
@@ 3.2.1 Discrete World Models via Prompting
Given a ToM problem $p$ and a constant $T\leq|p|$, we can rewrite $p$ as $p_{1}\oplus p_{2}\oplus\cdots\oplus p_{T}$. Our query function adds a standard preamble $x$ similar to that of CoT. DWM inserts, after each "split" $p_{t}$, an additional prompt $w$ like ‘Now, provide a succinct description of the state of the environment and each agent’s belief.’ and query an LLM to provide a representation of the current state description of the environment. An LLM is initially queried with $x\oplus p_{1}\oplus w$, and the answer $a_{1}$ is concatenated to the next query, i.e., $\psi(x\oplus p_{1}\oplus w\oplus a_{1}\oplus p_{2}\oplus w)$ to retrieve $a_{2}$ The process is carried on for each of the $T$ chunks, and, at the end, $y$ is concatenated to eventually prompt the model for the correct answer to $p$.
Let $z_{1}=\psi(x\oplus p_{1})$, $z_{t}=\psi(x\oplus p_{1}\oplus z_{1}\oplus p_{2}\oplus\dots\oplus z_{t-1}\oplus p_{t})=\psi(x\oplus\left(\bigoplus_{i=1}^{t-1}p_{i}\oplus z_{i}\right)\oplus p_{t})$, then, the final query is
$\psi(x\oplus\left(\bigoplus_{t=1}^{T}p_{t}\oplus z_{t}\right)\oplus y)$ (3)
In this sense, our partition function (Def. 3.2) consists of splitting a prompt into sequential chunks of the prompt, while the LLM is prompted to provide each state event at time $1\leq t<T$ as $e_{t}=\psi(x\oplus\left(\bigoplus_{t^{\prime}=1}^{t}p_{t^{\prime}}\oplus z_{t^{\prime}}\right)\oplus\omega)$. The process of prompting a model with DWM is illustrated in Figure 3.

### v3

@@ Abstract
Theory of Mind (ToM) can be used to assess the capabilities of Large Language Models (LLMs) in complex scenarios where social reasoning is required. While the research community has proposed many ToM benchmarks, their hardness varies greatly, and their complexity is not well defined. This work proposes a framework inspired by cognitive load theory to measure the complexity of ToM tasks. We quantify a problem’s complexity as the number of states necessary to solve it correctly. Our complexity measure also accounts for spurious states of a ToM problem designed to make it apparently harder. We use our method to assess the complexity of five widely adopted ToM benchmarks. On top of this framework, we design a prompting technique that augments the information available to a model with a description of how the environment changes with the agents’ interactions. We name this technique Discrete World Models (DWM) and show how it elicits superior performance on ToM tasks. [^1] Code and data for full reproducibility are available in the Code Material.
@@ 1 Introduction
Theory of Mind (ToM) studies how agents form and use beliefs to reason in dynamic environments Premack and Woodruff (1978). Originally developed to describe human interactions Preston and De Waal (2002); Tomasello (2009) as well as toddlers’ psychological development Wimmer and Perner (1983); Baron-Cohen et al. (1985), ToM has been quickly adopted by other fields, including artificial intelligence McCarthy (1979); Scassellati (2002), bayesian inference Baker et al. (2011) and machine learning Rabinowitz et al. (2018). In machine learning, ToM has both descriptive and prescriptive usage: on the one hand, ToM benchmarks assess the capabilities of a model in complex environments; on the other, ToM’s frameworks such as theory-theory Gopnik and Wellman (1994) and simulation theory Churchland (2013) have been widely adopted to test the proficiency of Large Language Models (LLMs) in social tasks where humans excel Strachan et al. (2024).
In this work, we propose a framework to characterise a ToM benchmark’s difficulty, i.e., its complexity, as the number of state events that are sufficient to track the state of an object, including $k^{\text{th}}$-order beliefs motivated by theoretical parallelisms with Sweller’s cognitive load theory Sweller (2010).
We characterise the complexity of five standard ToM benchmarks, from false belief to commonsense and social reasoning, and compute their complexity as a proxy of their inherent difficulty. Inspired by prompting techniques that split a task into elementary sub-problems that are solved sequentially, like Tree of Thoughts Yao et al. (2024) and least-to-most prompting Zhou et al. (2023a), we introduce a technique that stimulates a model’s reasoning capabilities via Discrete World Models (DWM). DWM leverages the notion of statefulness via a succinct and coherent representation of each state events, as illustrated in Figure 1. We test DWM on ToMi Le et al. (2019), MindGames Sileo and Lernould (2023), Adv-CSFB  Shapira et al. (2024), SocialIQA Sap et al. (2019), and FANToM Kim et al. (2023), eliciting superior performance than Chain of Thoughts (CoT) Wei et al. (2022) and Tree of Thoughts (ToT) Yao et al. (2024) on those problems whose state spaces are informative. We further assess whether memorisation affects a model’s performance, and we discover that while this phenomenon happens for standard benchmarks such as ToMi Le et al. (2019), with input-output pairs that can be retrieved word for word via prompting, it does not strongly correlate with a drop of performance on memorised ToM benchmarks. We conduct our experiments on a variety of open- and closed-source LLMs, including GPT-3.5-Turbo, GPT-4 OpenAI (2023), LLaMA3-70B AI@Meta (2024); Dubey et al. (2024) and Mixtral 8x7B Jiang et al. (2024). In summary, in this paper:
  • We introduce the concept of complexity of a ToM task to quantify the hardness of keeping track of the elements (e.g., agents’ beliefs or objects’ states) that are sufficient to produce the correct answer to different problems inspired by frameworks in cognitive science.
@@
We consider our work a step towards a framework that formalizes the hardness of a ToM problem in an unambiguous way, inspired by the theory of World Models Wong et al. (2023).
@@ 2 Related Work
Over 40 years of research on ToM in psychology Premack and Woodruff (1978); Baron-Cohen et al. (1985); Dennett (1988); Wellman (2017) on human development has created a fertile ground for the development of these ideas in adjacent fields. In the last decade, many works studied ToM in artificial intelligence and machine learning Baker et al. (2011); Rabinowitz et al. (2018), with applications to multi-agent systems and reinforcement learning Gronauer and Diepold (2022). More recently, the rise in popularity of LLMs shifted the interest towards understanding and benchmarking large models’ capacity to solve increasingly complex ToM tasks Aru et al. (2023); Zhou et al. (2023b); Mahowald et al. (2024). While some researchers believe LLMs have already become proficient in solving ToM tasks Bubeck et al. (2023); Kosinski (2023); Strachan et al. (2024), others show scepticism and illustrate cases where they fail on trivial variations of well-known problems Ullman (2023); Shapira et al. (2024); Sap et al. (2022). In a joint effort between computer scientists and psychologists, many ToM benchmarks have been developed and used to test neural-network models, including LLMs Gandhi et al. (2022); Chen et al. (2024); Strachan et al. (2024). Recently, concepts such as World Models Ha and Schmidhuber (2018) have found applicability mostly as discrete prompting techniques in conjunction with optimisation procedures Hao et al. (2023); Moghaddam and Honey (2023). Researchers have found evidence of an emergent internal representation (e.g., World Model’s surrogates) of the state games Li et al. (2023); Toshniwal et al. (2021) and state-tracking abilities Li et al. (2021); Kim and Schuster (2023); Kim et al. (2024), necessary for correct belief tracking in ToM problems. Cognitive load theory emerged in the late eighties with Sweller’s work on human problem solving Sweller (1988). Most measures of cognitive load are based on subjective reports from humans Sweller et al. (2011). Even though some attempts at automatic cognitive load measures have been present Yin et al. (2007), they have not been widely adopted in the community. The works that are more similar to our complexity framework are only tangentially related to ToM. Inspired by the work in Zhou et al. (2023a) and the results in Zhou et al. (2023b), our prompting technique is inspired by Park et al. (2023) and Nye et al. (2021): the former develops an architecture to record the agent’s experiences. The latter proposes a prompting technique that forces a model to express the intermediate computational steps to solve a problem.
Figure 2: How statefulness and statelessness (Def. 3.2) are computed for the motivating example in Fig. 1. For ${obj_{1}}$, an optimal split to track the apple merges the first two states and chunks of the input prompt. For ${obj_{2}}$, which involves the $1^{\text{st}}$-order belief of Bob, the statefulness is higher, with $e_{2}$ that cannot be merged with $e_{3}$ as it introduces partial observability. The complexity of the task (bottom) is computed as per Eq. 2, where the complexity of objects that are not directly relevant to the question/answer is discounted.
@@ 3 Methodology
In this section, we introduce a notion of complexity for ToM problems which quantifies the hardness of a problem as the number of computational steps humans take to solve them and compare it with Sweller’s cognitive load theory. We then present the DWM prompting technique within the complexity framework and show how it differs from standard methods like CoT and ToT. We further characterise its efficiency with the number of input/output tokens and queries to a model as the control variables.
@@ 3.1 On the Complexity of ToM
The need to provide a consistent representation of the environment, including each agent’s beliefs, inspired us to characterise the complexity of a ToM problem in terms of sufficient elements to track to output the correct result. Consider a problem prompt $p$, expressed in natural language, that describes how multiple agents interact with an environment object $\mathbf{obj}$, as illustrated in Figure 2 (top). In our framework, an object can be the state of the apple as well as the $k^{\text{th}}$-order belief of an agent about the apple position. Our framework naturally extends to multiple objects by considering their union.
Suppose that in $p$, the state of $\mathbf{obj}$ is modified $T>0$ times, thus identifying $T$ unique configurations, namely $E_{\mathbf{obj}}=\{e_{1},..,e_{T}\}$. To correctly solve a ToM task where $p$ is complemented by a query about $\mathbf{obj}$, a model should distinguish between the interactions that modify the configuration of $\mathbf{obj}$, i.e., the stateful states, from those that modify any other stateless object $Obj\setminus\mathbf{obj}$, i.e., those that one does not need to track.
We first show how to define the cost of tracking a task’s stateful states, which we complement with that of the stateless. Both definitions concur in defining the complexity of a ToM task.
@@ 3.1.1 Stateful and Stateless Complexity
A state event for an object $\mathbf{obj}$ is an event that links adjacent state descriptions that involve, for both the environment state $e_{t}$ and the sub-prompt $p_{\leq t}$, a state change of $\mathbf{obj}$. Formally, we define a relation, $F_{\mathbf{obj}}$, to specify which pairs of state descriptions form a state event: $F_{\mathbf{obj}}((e_{t},p_{\leq t}),(e_{t+1},p_{\leq t+1}))\equiv e_{t}\neq e_{t+1}\ \wedge\ p_{\leq t+1}=p_{\leq t}\oplus p_{t+1}$ where $1\leq t\leq\lvert p\rvert.$ ($\lvert p\rvert$ denotes the number of atomic prompts.) and $\oplus$ is the string concatenation operator.
@@ 3.1.2 The Complexity of a ToM Task
We can now define the notion of statefulness of a ToM task specified as a prompt $p$ as the size of Eq. 3.2, namely $T_{\mathbf{obj}}=\lvert part_{\mathbf{obj}}\rvert$. The process of computing the statefulness of an object or its belief is illustrated in Fig. 2.
For a ToM task where the question to solve relates to an object $\mathbf{obj}$, one must ensure that changes to any other object, namely $Obj\setminus\mathbf{obj}$, do not affect $\mathbf{obj}$. While tracking the evolution of what is irrelevant to answer the question is unnecessary, a computation model must assess whether a particular environmental change affected $\mathbf{obj}$. We thus introduce the notion of statelessness, i.e., the cost of discerning whether a change in the environment affects $\mathbf{obj}$. The computation is similar to that of Def. 3.2, except that $\mathbf{obj}$ is replaced with any object in $Obj\setminus\mathbf{obj}$; however for stateless objects, we introduce a discount factor $\tau$ to penalise the complexity of state events that do not affect $\mathbf{obj}$. Mathematically, we formalise the statelessness of a ToM task involving an object $\mathbf{obj}$ as $\tau\sum_{obj\in Obj\setminus\mathbf{obj}}T_{obj}$.
@@
3.1.3 A parallelism from the Cognitive sciences
Understanding how humans solve complex problems has long served as a valuable source of inspiration for advancing machine intelligence. Thought frameworks in the cognitive sciences, such as Kahneman’s Dual-process theory Kahneman (2011), have greatly influenced various fields, including artificial intelligence. In this work, we draw ideas from another theory, less known in the community, the cognitive load theory Sweller (1994, 1988, 2010). According to this theory, three main factors influence the mental effort humans exert when solving a particular task or learning new information: intrinsic, extraneous and germane load. The intrinsic load measures the natural difficulty of a certain task, the information that needs to be digested before answering the question. This relates to the complexity of the material itself. The extraneous load concerns the presentation of the information in the problem. For example, if questions phrased in a complex manner are used with a child, it would be much more difficult to understand and answer correctly compared to easier phrasing. Similarly, if many confounding sentences or sentences that do not matter in answering a question are present in the text, we expect an LLM to be worse, suggesting a weak similarity between the means of reason of these models and humans. Finally, germane load estimates the working memory resources needed to understand the important parts of the problem, i.e. the intrinsic load. If part of the memory is devoted to the extraneous load, then the germane load is diminished, suggesting a positive correlation with intrinsic load and a negative correlation with extraneous load.
3.1.4 A comparison with the Cognitive Load Theory
Our framework, summarized in figure 2, has two main parts: stateful and stateless complexity. These notions have some similarities with, respectively, the intrinsic load and extraneous load. Stateful complexity provides a measure on the sentences that are needed to answer the question correctly and must be adequately represented in memory. In a similar manner, intrinsic load concerns on the needed information to correctly analyze a task. Likewise stateless complexity yields information about the confounding or irrelevant sentences and phrases in the text akin to extraneous load. In our setting, germane load could be interpreted as the ratio of the stateful and stateless complexity: higher ratio means higher density of useful sentences in answering a question. This notion of load could be used as a basis of an objective measure on the quality of a question-answering sample: given the same quantity of cognitive load, i.e. complexity, we would like to have a simple presentation with correct information, maximizing the germane load. If the cognitive load hypothesis applies to LLMs, maximizing the germane load would lower the complexity of the tasks given to a model, and therefore it would aid the model to answer questions more accurately.
@@ 3.2 Discrete World Models
A (Large) Language Model is a function that predicts the next token (out of a finite vocabulary) conditioned on the sequence of previously fed/generated tokens, namely $\psi:\mathbf{v}\in V^{*}\xrightarrow{}v\in V$. Such a mechanism can be used to sample multiple token outputs until an ‘end-of-text’ token is predicted, by invoking $\psi$ in an auto-regressive fashion, i.e., $\psi(v|\mathbf{v})$. In our setting, a problem is specified as a tuple $(p,Q)$, where $p$ is a ToM problem and $Q$ is a query function that modifies $p$ according to a prompting technique, namely $Q:p\xrightarrow{}p^{\prime}$. The LLM’s output $y$ for an input $Q(p)$ is then compared for correctness against an oracle $\Omega$, i.e., $\Omega:\psi(Q(p))\xrightarrow{}\{0,1\}$, where $1$ means correct classification (00, otherwise). On a sample of $N>0$ ToM problems, the accuracy of a model $\psi$ is then measured as $\frac{1}{N}\sum_{i=1}^{N}\Omega(\psi(Q(p_{i}))$, i.e., the average number of times a model is correct in its prediction.
@@ 3.2.1 Discrete World Models via Prompting
Given a ToM problem $p$ and a constant $T\leq|p|$, where $|p|$ is ideally measured as the number of state changes in the problem, we can rewrite $p$ as $p_{1}\oplus p_{2}\oplus\cdots\oplus p_{T}$. Our query function adds a standard preamble $x$ similar to that of CoT. DWM inserts, after each “split” $p_{t}$, an additional prompt $w$ like ‘Now, provide a succinct description of the state of the environment and each agent’s belief.’ and query an LLM to provide a representation of the current state description of the environment. An LLM is initially queried with $x\oplus p_{1}\oplus w$, and the answer $a_{1}$ is concatenated to the next query, i.e., $\psi(x\oplus p_{1}\oplus w\oplus a_{1}\oplus p_{2}\oplus w)$ to retrieve $a_{2}$ The process is carried on for each of the $T$ chunks, and, at the end, $y$ is concatenated to eventually prompt the model for the correct answer to $p$.
Let $a_{1}=\psi(x\oplus p_{1}\oplus w)$, $a_{t}=\psi(x\oplus p_{1}\oplus w\oplus a_{1}\oplus p_{2}\oplus\dots\oplus a_{t-1}\oplus p_{t})=\psi(x\oplus\left(\bigoplus_{i=1}^{t-1}p_{i}\oplus w\oplus a_{i}\right)\oplus p_{t})$, then, the final query is
$\psi(x\oplus\left(\bigoplus_{t=1}^{T}p_{t}\oplus w\oplus a_{t}\right)\oplus y)$ (3)
In this sense, our partition function (Def. 3.2) consists of splitting a prompt into sequential chunks of the prompt, while the LLM is prompted to provide each state event at time $1\leq t<T$ as $e_{t}=\psi(x\oplus\left(\bigoplus_{i=1}^{t}p_{i}\oplus w\oplus a_{i}\right)\oplus\omega)$. The process of prompting a model with DWM is illustrated in Figure 3.